# RTop-K: Ultra-Fast Row-Wise Top-K Algorithm and GPU Implementation for Neural Networks

### v1

@@ Introduction
The performance of RTop-K with different early stopping settings is compared against the row-wise top-$k$ implementation provided by PyTorch (Pytorch 2024), which is the state-of-the-art row-wise top-$k$ implementation that can support a large number of rows. The latency measurements are conducted using the Nsight Compute (NVIDIA 2023) tool.
Result Analytics
@@
Fig. 5 presents a comprehensive evaluation of RTop-K compared to the PyTorch implementation. It can be observed that RTop-K demonstrates significant speed improvements over PyTorch across various early stopping $max\_iter$ settings. Even with no early stopping ($\epsilon=10^{-4}$), RTop-K consistently outperforms PyTorch in all scenarios. Moreover, the speed of RTop-K remains nearly unaffected by different values of $k$. Table 3 summarizes the average speed-up of RTop-K relative to PyTorch, showing a speed increase of 4.245×4.245$\times$4.245 × to 9.506×9.506$\times$9.506 × with early stopping, and 3.936×3.936$\times$3.936 × with no early stopping.
Table 4 summarizes the accuracy of four real-world MaxK-GNN based GraphSage models and the proportion of time spent on row-wise top-k operations during training. It is evident that row-wise top-k operations account for a substantial portion of the training time, ranging from 11.66% in Reddit to 26.86% in Flickr. This indicates that optimizing top-k operations in real-world MaxK-GNN based models is meaningful for their training. The impact of applying RTop-K with different early stopping settings in the actual training of these four models is shown in Fig. 4. It can be observed that the testing accuracy of the models remains good, and in the cases of the Reddit, Flickr, and Yelp models, the testing accuracy of RTop-K with early stopping even surpasses that of the optimal top-k. This demonstrates the successful effectiveness of RTop-K in real-world applications.
@@
Table 3: Average speed up of RTop-K compared to PyTorch implementation ($\epsilon=10^{-4}$ for no early stopping).
Max Iter Avg Speed Up Max Iter Avg Speed Up
2 9.506 6 5.256
3 8.216 7 4.699
4 6.965 8 4.245
5 6.002 no early stopping 3.936
@@
Table 4: Graph data and the testing accuracy/F1 score of the MaxK-GNN based GraphSage model along with the percentage of time spent on row-wise top-k operations during training.
Graph # Nodes Acc/F1 score (%) Top-k time (%)
ogbn-products 2449029 80.59 19.81
yelp 716847 61 26.09
reddit 232965 96.65 11.66
flickr 89250 53.6 26.86
Conclusion
@@
In this paper, we presented RTop-K, a highly efficient parallel row-wise top-k selection algorithm for GPUs. By employing a Binary Search-based approach, RTop-K significantly accelerates top-k operations while maintaining the accuracy of neural network models, as confirmed by our theoretical analysis. Comprehensive testing showed that RTop-K outperforms state-of-the-art GPU implementations, achieving speed-ups of 4.245× to 9.506× with early stopping and 3.936× without early stopping. These findings demonstrate RTop-K’s potential to improve the performance of Graph Neural Networks (GNNs) by addressing challenges in latency and throughput on GPU platforms.

### v3

@@ 4.1 Setup and Configuration
The CUDA source code of RTop-K is compiled using NVCC, version 12.6, and executed on an NVIDIA A6000 platform running Ubuntu 22.04. We conduct comprehensive performance tests on the RTop-K kernel, covering various input matrix dimensions, with the number of rows $N$ ranging from $2^{14}$ to $2^{20}$, hidden dimensions $M$ ranging from 256 to 768, and $k$ values ranging from 16 to 128. In all cases, we evaluate the speed of RTop-K with different early stopping settings, including $max\_iter$ values from 2 to 8, as well as no early stopping ($\epsilon=10^{-16}$). The results are compared against the row-wise top-$k$ implementation provided by PyTorch (Pytorch, 2024), which is the state-of-the-art row-wise top-$k$ implementation supporting a large number of rows. The latency measurements are conducted using the Nsight Compute tool (NVIDIA, 2023).
We also integrate the RTop-K kernel into MaxK-GNN models to evaluate the overall speedup of the entire training process and the impact of different early stopping settings on test accuracy. The evaluation covers three models in MaxK-GNN: GraphSAGE (Hamilton et al., 2017), GCN (Kipf & Welling, 2016), and GIN (Xu et al., 2019). The graph datasets used include Flickr (McAuley & Leskovec, 2012), Yelp (Zeng et al., 2020), Reddit (William L. Hamilton, 2017), and Ogbn-products (Hu et al., 2020).
4.2 RTop-K Kernel Evaluation
@@
Fig. 4 presents a comprehensive time profiling of RTop-K compared to the PyTorch implementation. It can be observed that RTop-K demonstrates significant speed improvements over PyTorch across various early stopping $max\_iter$ settings. Even with no early stopping, RTop-K significantly outperforms PyTorch in all scenarios. Moreover, the dimension that has the most impact on the speed-up ratio is $M$, while $N$ and $k$ have relatively smaller effects. Table 3 summarizes the average speed-up of RTop-K relative to PyTorch for different values of $M$. When $M=256$, the speed-up varies from 9.55×9.$5$ to 13.07×13.07$\times$13.07 × across different $max\_iter$ settings, and even with no early stopping, the speed-up is still as high as $8\times8$. On average, the speed-up ranges from 7.99×7.$9$ to 11.49×11.49$\times$11.49 × with early stopping, and 7.29×7.29$\times$7.29 × with no early stopping.
We observe that the speed-up with no early stopping is close to that of $max\_iter=8$, which indicates that although the binary search requires many iterations in the worst-case, it typically exits early in most cases. This observation is consistent with the results presented in Table 1. Even with a few bad cases, the overall kernel speed remains unaffected.
@@
Table 3: Average speed up of RTop-K compared to PyTorch implementation ($\epsilon=10^{-16}$ for No Early Stopping) across different $M$ values.
Max Iteration 2 3 4 5 6 7 8 No Early Stopping
$M$=256 13.07 12.32 11.46 10.86 10.32 9.88 9.55 8.88
$M$=512 11.66 11.37 10.43 9.51 8.87 8.34 7.98 7.27
$M$=768 9.73 9.44 8.72 7.75 7.16 6.78 6.46 5.72
Average 11.49 11.04 10.20 9.37 8.79 8.34 7.99 7.29
@@
Table 4: Graph data and the baseline testing accuracy of the MaxK-GNN based GNN model along with the percentage of time spent on row-wise top-k operations during training.
GNN Model GraphSAGE GCN GIN
Graph #Nodes Acc(%) Top-k Prop(%) Acc(%) Top-k Prop(%) Acc(%) Top-k Prop(%)
Ogbn-products 2449029 80.08 19.81 76.6 19.61 77.77 19.67
Yelp 716847 61.09 26.09 48.26 25.84 43.16 25.92
Reddit 232965 96.74 11.66 95.18 11.61 94.96 11.62
Flickr 89250 53.44 26.86 50.42 26.78 51.73 26.73
4.3 Model Training and Testing Performance Evaluation
Table 4 summarizes the proportion of time spent on row-wise top-k operations in several MaxK-GNN training instances. It is evident that row-wise top-k operations account for a substantial portion of the training time, ranging from 11.61% in the Reddit GCN training to 26.86% in the Flickr GraphSAGE training. This indicates that optimizing row-wise top-k operations is meaningful for improving their training efficiency.
The impact of applying RTop-K on speed and accuracy with different early stopping settings in the actual training of these models is shown in Fig. 5, with the setting of $N=\#\text{Nodes}$, $M=256$, and $k=32$. For all GNN models and all graphs, the application of RTop-K effectively accelerates the overall training workflow. Specifically, under different $max\_iter$ settings, the average overall training speed-up for Reddit ranges from 11.97% to 12.21%, for Flickr from 32.48% to 33.29%, for Ogbn-products from 22.00% to 22.74%, and for Yelp from 31.21% to 32.42%.
[…]
It can be observed that the testing accuracy of the models remains high. Except for the GIN training on Flickr, the testing accuracy across different $max\_iter$ settings for other cases fluctuates around the testing accuracy achieved with the optimal row-wise top-k selection. In many cases, applying early stopping for row-wise top-k selection even results in higher testing accuracy. This is also a manifestation of the inherent robustness of GNNs.
5 Conclusion
@@
In this paper, we presented RTop-K, a highly efficient parallel row-wise top-k selection algorithm for GPUs. By employing a binary search-based approach, RTop-K significantly accelerates top-k operations while maintaining the accuracy of neural network models, as confirmed by our theoretical analysis. Comprehensive kernel evaluations demonstrate that RTop-K outperforms state-of-the-art GPU implementations, achieving an average speed-up of up to 11.49×11.49$\times$11.49 × with early stopping and 7.29×7.29$\times$7.29 × without early stopping. The evaluation of the overall MaxK-GNN training workflow with RTop-K shows that RTop-K provides overall speed-ups ranging from 11.97% to 33.29% across different models and datasets, with early stopping having almost no impact on testing accuracy.
6 Acknowledgments
This research was supported in part by NSF SHF-2505770, Semiconductor Research Corporation (SRC) Artificial Intelligence Hardware program.
Appendix A The expectation of the iteration counts of Algorithm 1
Assume that a vector $\mathbf{v}$ of length $M$ has elements that follow a normal distribution $N(\mu,\sigma^{2})$. The probability that an element $x$ exceeds a threshold $thres$ is given by $P(x>thres)=1-\Phi\left(\frac{thres-\mu}{\sigma}\right)$, where $\Phi$ is the cumulative distribution function of the normal distribution. If the expected number of elements selected from $\mathbf{v}$ is $k$, the expectation of $thres$, denoted as $E(thres)$, satisfies:
$M\cdot\left(1-\Phi\left(\frac{E(thres)-\mu}{\sigma}\right)\right)=k\implies E(thres)=\mu+\sigma\cdot\Phi^{-1}\left(1-\frac{k}{M}\right)$ (1)
Considering the distinguishable interval $\delta$ between the $k$-th and $(k+1)$-th largest elements, the length of this interval is:
$\delta=\frac{1}{M\cdot f(E(thres))}$ (2)
where
$f(E(thres))=\frac{1}{\sigma\sqrt{2\pi}}\exp\left(-\frac{(E(thres)-\mu)^{2}}{2\sigma^{2}}\right)$
is the probability density at $E(thres)$. The length of the initial search interval $D$ is given by:
$D=\text{max}(\mathbf{v})-\text{min}(\mathbf{v})\approx 2\sigma\sqrt{2\ln M}$ (3)
Each iteration of binary search halves the search interval length and moves closer to $E(thres)$. The expected number of iterations $E(n)$ required for the algorithm to exit is determined by the search interval shrinking to within $\delta$. Thus, $E(n)$ can be approximated as:
$E(n)$ $\approx\log_{2}\left(\frac{D}{\delta}\right)=\log_{2}\left(2\sigma\sqrt{2\ln M}\cdot M\cdot f(E(thres))\right)$
$=\log_{2}\left(2M\sqrt{\frac{\ln M}{\pi}}\right)-\frac{1}{2\ln 2}\left(\Phi^{-1}\left(1-\frac{k}{M}\right)\right)^{2}$ (4)
Table 5: Cumulative percentage of iterations where the loop exits in Algorithm 1, for different $M,k$ values, with $\epsilon=0$. Experimental results are based on $10^{4}$ repeated experiments for each $M,k$ couple, and theoretical values $E(n)$ are also provided.
Iters M, k
256
64
256
128
1024
64
1024
128
1024
256
1024
512
4096
64
4096
128
4096
256
4096
512
8192
64
8192
128
8192
256
8192
512
1 0.12 1.5 0 0 0.02 0.53 0 0 0 0 0 0 0 0
2 0.17 1.5 1.25 0.17 0.02 0.53 0.41 0.41 0.18 0 0.12 0.23 0.24 0.02
3 1.98 1.62 1.27 0.5 0.62 0.54 0.43 0.41 0.2 0.08 0.3 0.24 0.25 0.02
4 3.48 2.9 1.93 1.31 0.99 0.7 1.54 0.63 0.36 0.33 1.11 0.7 0.32 0.15
5 6.88 6.06 4.05 2.42 1.67 1.25 2.83 1.34 0.76 0.61 2.43 1.3 0.7 0.31
6 13.96 11.64 8.18 4.49 3.24 2.64 5.4 2.84 1.64 1.06 4.68 2.62 1.46 0.75
7 27.3 23.03 16.09 9.42 6.41 5.16 11.13 5.99 3.37 2.15 9.18 5.28 2.98 1.6
8 46.47 40.18 30.63 18.77 12.7 10.19 21.56 12.05 6.89 4.15 19.25 10.62 5.59 3.27
9 66.34 60.57 50.71 35.21 24.91 19.69 39.04 24.37 13.69 8.23 34.8 21.5 11.49 6.72
10 81.07 76.76 69.08 55.43 43.92 35.98 60.23 42.59 26.65 16.71 55.41 38.13 22.42 13.3
11 90.11 87.67 82.99 72.93 63.57 55.7 76.59 62.55 46.7 31.29 73.09 58.45 39.96 25.94
12 95.17 93.48 90.71 84.79 78.45 73.32 87.42 78.11 66.36 52.28 85.21 75.11 60.66 44.97
13 97.46 96.8 95.19 91.9 88.37 85.06 93.42 87.9 81.06 70.87 92.16 86.37 77.12 64.65
14 98.68 98.35 97.59 95.86 93.82 92.47 96.48 93.18 89.91 83.91 96.29 92.63 87.55 80.25
15 99.46 99.21 98.79 97.95 96.87 96.32 98.24 96.3 94.73 91.34 98.15 96.39 93.48 89.3
16 99.73 99.56 99.43 98.89 98.46 98.06 99.21 98.14 97.35 95.39 99.11 98.19 96.73 94.46
17 99.85 99.72 99.71 99.51 99.18 99.03 99.59 99.02 98.51 97.65 99.48 99.18 98.35 97.29
18 99.96 99.85 99.9 99.73 99.62 99.44 99.77 99.45 99.2 98.8 99.77 99.59 99.21 98.66
19 99.99 99.93 99.95 99.84 99.84 99.67 99.87 99.75 99.64 99.38 99.89 99.76 99.54 99.33
20 99.99 99.98 99.97 99.9 99.93 99.85 99.95 99.87 99.84 99.76 99.98 99.89 99.8 99.69
21 99.99 99.99 99.98 99.95 99.97 99.93 99.97 99.94 99.95 99.95 99.99 99.98 99.94 99.84
22 99.99 99.99 99.99 99.96 99.99 99.95 99.99 99.98 99.98 99.97 99.99 99.98 99.96 99.9
23 99.99 99.99 99.99 99.99 99.99 99.96 100 99.98 100 99.98 100 100 99.99 99.94
24 100 99.99 100 100 99.99 99.97 - 100 - $9$ 100 99.96
25 - $9$ 100 $9$ $9$ 99.98
26 - 100 - - - 100 - - - $9$ 99.99
27 - - - - - - - - - 100 - - - 99.99
28 - - - - - - - - - - - - - 100
Avg 8.72 9 9.53 10.31 10.87 11.24 10.07 10.95 11.73 12.46 10.3 11.14 12.02 12.8
$E(n)$ 9.08 9.41 9.87 10.62 11.24 11.57 10.36 11.2 12 12.75 10.54 11.41 12.26 13.06
We compared the calculation results of Equation (4) with more detailed experimental results, as shown in Table 5. It can be observed that the results match well, but $E(n)$ is always slightly larger than the measured average exit. This could be because the estimation of the initial search interval, $D\approx 2\sigma\sqrt{2\ln M}$, is valid only when $M$ is sufficiently large. When $M$ is not large enough, the lack of tail samples causes the actual initial search interval to be smaller.
Appendix B A comprehensive analysis of the performance of RTop-K when applied to varying vector sizes
Our design fixes one warp to process one vector. As the vector size $M$ increases, the shared memory required per warp also increases. Given that the available shared memory per block has a limit, on the A6000 GPU, we allocate only $\lfloor 8192/M\rfloor$ warps per block. For $M>8192$, our current shared memory-based acceleration strategy cannot be directly applied.
For $M\leq 8192$, the speedup of RTop-K relative to PyTorch is shown in Figure 6.
Figure 6: Speedup of RTop-K relative to PyTorch for different vector sizes $M$ and number of vectors $N=65536$. The average speedup for each case is computed over $k=64,128,256,512$ and $k<M$. Precision $\epsilon=10^{-16}$ is used for the no early stopping version.
Considering the lower-bound speed of RTop-K (no early stopping version):
  • When $M$ is below 1280, RTop-K achieves more than a 5×5$\times$5 × speedup over PyTorch.
  • When $M$ is between 1280 and 3072, RTop-K achieves a 2×2$\times$2 × to 5×5$\times$5 × speedup over PyTorch.
  • When $M$ is between 3072 and 6144, RTop-K achieves a 1×1$\times$1 × to 2×2$\times$2 × speedup over PyTorch.
  • When $M$ is between 6144 and 8192, RTop-K is slower than PyTorch, with only the early stopping version using a small $max\_iter$ still being faster than PyTorch.
Theoretically, as shown in Equation (4), the expected number of search iterations for Algorithm 1 is:
$E(n)=\log_{2}\left(2M\sqrt{\frac{\ln M}{\pi}}\right)-\frac{1}{2\ln 2}\left(\Phi^{-1}\left(1-\frac{k}{M}\right)\right)^{2}<\log_{2}\left(2M\sqrt{\frac{\ln M}{\pi}}\right)=O(\log M)$
In each iteration, a reduction operation of length $M$ is required. Since one warp is fixed for one vector, the time complexity of this reduction is equivalent to serial reduction, which is $O(M)$. Therefore, the total time complexity of Algorithm 1 is $O(M\log M)$. In contrast, PyTorch’s underlying operation, RadixSelect, has a time complexity of $O(M)$. Thus, when $M$ is sufficiently large, Algorithm 1 will lag behind traditional algorithms.
However, from a practical perspective, for $M\leq 8192$, as shown in Table 5, the growth of $E(n)$ is very slow. Additionally, since the search range has a lower bound (depending on the data type), the number of search iterations has an upper bound. Moreover, the searching stage is fully executed in shared memory, which leads to a decreasing proportion of time spent in the searching stage compared to the loading and selecting stages, as these involve increasing global memory accesses with the growth of $M$.
Therefore, we believe the actual time complexity of Algorithm 1 is less than $O(M\log M)$. As $M$ increases, the relative speedup of RTop-K compared to PyTorch decreases primarily because the relative efficiency of PyTorch’s RadixSelect improves (the proportion of its initialization, histogram construction, and indexing overhead decreases).
We also evaluated the performance of Algorithm 1 with different precision settings, and the results are shown in Figure 7.
Figure 7: Speedup of RTop-K (no early stopping version) relative to PyTorch for different vector sizes $M$ and different precisions, with the number of vectors $N=65536$. The average speedup for each case is computed over $k=64,128,256,512$ and $k<M$.
We found that precision has almost no impact on speed. Even with the setting of $\epsilon=0$, as shown in Table 5, Algorithm 1 exits within 16 iterations in the vast majority of cases, and the remaining rare cases have a negligible impact on the overall performance. This is also because the searching stage is fully executed in shared memory, making Algorithm 1 less sensitive to the number of search iterations.